Exercise 3.3 A web page designer creates an animation in which a dot on a computer screen has a position of r42 cm(2.1 cm/s2 )ti (4.7 cm/s )tj Part A Find the average velocity of the dot between t 0 and t- 2.0 s Give your answer as a pair of components separated by a comma. For example, if you think the x component is 3 and the y component is 4, then you should enter 3,4. Express your answer using two significant figures ANSWER cm/s ate Part B Find the instantaneous velocity at t 0 Give your answer as a pair of components separated by a comma. For example, if you think the x component is 3 and the y component is 4, then you should enter 3,4. Express your answer using two significant figures ANSWER cm/s Part C Find the instantaneous velocity at t 1.0 s Give your answer as a pair of components separated by a comma. For example, if you think the x component is 3 and they component is 4, then you should enter 3,4. Express your answer using two significant figures ANSWER cm/s Part D Find the instantaneous velocity att 2.0s Give your answer as a pair of components separated by a comma. For example, if you think the x component is 3 and the y component is 4, then you should enter 3,4. Express your answer using two significant figures ANSWER cm/s
Answer #1

A) displacement at t=2;

r2 = (4.2 + 2.1*22)i + 4.7*2; r2 = 12.6i + 9.4j

displacement at t=0;

r0 = (4.2 + 2.1*02)i + 4.7*0; r2 = 4.2i

net dispacement

r = r2 - r0 = (8.4i + 9.4j ) cm

average velocity, = displacement / time = (8.4i + 9.4j )/ (2-0)

average velocity, = 4.2 i + 4.7 j

4.2, 4.7

function of velocity

v = 4.2*t i + 4,7 j

B) velocity at t=0

v0= 4.2*0 i + 4.7 j

0, 4.7

C) velocity at t=1

v1= 4.2*1 i + 4.7 j

4.2, 4.7

D) velocity at t=2

v1= 4.2*2 i + 4.7 j

8.4, 4.7